OpenStudy (anonymous):

Solve this ∫▒〖cos^2⁡x dx〗 = tanx + c

Parth (parthkohli):

\[\textbf{HINT}\]Use \(\cos^2 (x) = \dfrac{\cos(2x)+1}{2}.\)

OpenStudy (anonymous):

sorry i've wrote wrong exercise it's \[\int\limits_{?}^{?}1/\cos^2x dx = tanx + c \]